a piggy bank contains 30 coins, all dimes and quarters worth 5.70. how many dimes and quarters are in the bank?
Answer by tommyt3rd(5050)

D+Q=30
.1D+.25Q=5.7 (m by 20)
D+Q=30
2D+5Q=114

D+Q=30 (m by -2)
2D+5Q=114
-2D-2Q=-60
2D+5Q=114

adding...
3Q=54
Q=18
D=30-18=12
12 dimes, 18 quarters